Adjusted WACC. Hollydale's is a clothing store in East Park. It paid an annual dividend of ​$2.60 last year to its shareholders and plans to increase the dividend annually at 4.5​%. It has 540,000 shares outstanding. The shares currently sell for $28.06 per share. ​ Hollydale's has 14,000 semiannual bonds outstanding with a coupon rate of 6.5​%, a maturity of 19 years, and a par value of ​$1,000. The bonds are currently selling for ​$591.73 per bond. What is the adjusted WACC for​ Hollydale's if the corporate tax rate is 20​%?

What is the adjusted WACC for​ Hollydale's if the corporate tax rate is 20​%?

​(Round to two decimal​ places.)
Answer #1

To compute WACC, we need to arrive at the Cost of Debt and Cost of Equity

Cost of Equity will be arrived at using the Gordon Growth Model, which states that

Market Price of Stock = Current Dividend * (1 + Growth Rate) / (Required Rate of Return on Equity - Growth Rate)

Market Price of Stock = $28.06

Current Dividend = $2.60

Growth Rate = 4.5%

Hence, plugging in the respective values into the equation we get

Required Rate of Return on Equity = 14.18%

Hence, cost of equity = 14.18%

Cost of debt = 6.5% * (1-0.20) = 5.2%

Market value of Equity = (540,000 * 28.06) = 15,152,400

Market value of Debt = (14,000 * 591.73 ) = 8,284,220

Weight of equity = 64.6%

Weight of Debt = 35.4%

WACC= (0.646 * 14.18%) + (0.354 * 5.2%) = 11.01%

WACC = 11.01%